Suppose that it snows in Greenland an average of once every 21 days, and when it does, glaciers have a 26% chance of growing. When it does not snow in Greenland, glaciers have only a 9% chance of growing. What is the probability that it is snowing in Greenland when glaciers are growing?
Answer #1

P(Snow) = 1/21 = 0.0476

So P(Not snow) = 1 - 0.0476 = 0.9524

P(Glacier grow | Snow) = 0.26

P(Glacier grow | Not snow) = 0.09

Hence by Bayes' theorem:

P(Snow | Glacier grow)

10.26 + 0.0476 -0.26 * 0.0476 +0.00 * 0.9524

= 0.1262